# A Neural Rewriting System to Solve Algorithmic Problems

### v1

@@ 4.2 Solver Module
Differently from classical rewriting systems, we do not represent rewriting rules symbolically in a dictionary of rules. Instead, we aim to learn rewriting rules from data using a neural network. Given a leaf formula $f^{L}$ generated by the Selector, the Solver is trained to output the corresponding reduction $e$, thereby implementing an appropriate rewriting rule for the formula. When training the Solver, we assume that the Selector will always produce perfect outputs (teacher forcing). The Solver is also trained to identify the termination state of the computation. That is, when the formula has reached its final state, it is the Solver’s responsibility to signal that such state has been reached. The Solver thus also learns to map atomic elements, which correspond to the final value of a formula, to the special symbol $\omega$ that represents the end of computation.
As in the case of the Selector, we frame the task as a sequence-to-sequence task. In this case, since the Solver has to compute simple mappings without any particular assumption, we implement it using a vanilla transformer.
@@ 4.3 Combiner Module
The last module in the architecture is the Combiner, whose goal is to produce a simplified version of the original formula, given the formula itself $f$, the leaf formula $f^{L}$ identified by the Selector and its reduction $e$ computed by the Solver.
@@
More precisely, we represent both the input sequence $f$ and the leaf formula $f^{L}$ as sequences of 1-hot vectors over the same vocabulary. Since the leaf formulas found for different sequences in a batch can have different lengths, we pad each one with zeros to prevent the padding to match in the input. Then, we set the filter of the 2D CNN to the 1-hot representation of $f^{L}$. We refer to the CNN parameterized in this way as $\mathrm{CNN}_{f^{L}}$. Doing so allows us to obtain from the output of the convolution both information on the location of the best match of $f^{L}$ in $f$ and on the number of tokens in $f^{L}$ that match $f$ exactly in some point. Indeed, we can compute the location of the best match as $\mathrm{pos}(f^{L},f)=\mathrm{argmax}~{}\mathrm{CNN}_{f^{L}}(f)$. Furthermore, we can compute the agreement score $a(f^{L},f)=\frac{\mathrm{max}~{}\mathrm{CNN}_{f^{L}}(f)}{|f^{L}|}$, where $|f^{L}|$ is the number of tokens in $f^{L}$. Dividing by $|f^{L}|$ makes the score normalized, which allows us to compare the agreement scores of leaf formulas with different lengths. Indeed, as described in Section 4, the Selector uses this score in the multi-output generation strategy to discard the outputs that do not have an exact match in the input formula. Notice that in this case the CNN is parameterized using candidate leaf formulas $\hat{f}^{L}_{j}$ whose accuracy scores with $f$ are compared. If there is no Selector output such that $a(\hat{f}^{L}_{j},f)=1$, the computation on the input sequence $f$ is interrupted, and this is considered as a failure of the model.
@@ 5 Experiments and results
(a) ListOps
(b) Arithmetic
(c) Algebra
Figure 3: Performance of the Neural Data Router, GPT-4 and Neural Rewriting System on the ListOps, Arithmetic and Algebra problems.
We evaluated the capacity of our model to solve out-of-distribution problem instances by considering three cases of formula simplification tasks: ListOps, arithmetic, and algebraic formulas. We benchmark our Neural Rewriting System against the Neural Data Router and GPT-4, which is one of the most advanced LLMs currently available. Code and data used in our experiments are available as Supplementary Materials and will be made public upon acceptance.
@@ 5.1 Datasets
For all three tasks, the automatic generation of formulas is parameterized by specifying the nesting level of a formula, that is the maximum number of nested formulas that can be found as arguments of the top-level formula plus the level of the formula itself. At each level, any formula is then nested in two points: exactly two arguments in the formulas on that level will be other formulas.
ListOps The ListOps dataset Nangia and Bowman (2018) was introduced to evaluate the capacity of neural networks to build parse trees of nested formulas. The original dataset consists of formulas composed of operations on lists of integers, including minimum, maximum, median and sum modulo $10$ of a list of integers. We modified the ListOps dataset to generate expression that comply with the schema described above. In particular, we fixed the number of nesting points at each level of a formula to 2, and we made it possible to specify the number of arguments that appear in formulas at any level. To reduce the complexity of the problem, we reduced the set of operations to minimum, maximum and sum modulo $10$.
@@
Algebra We considered a subset of algebraic formulas which can always be reduced up to a minimal form in a deterministic way. Such formulas involve sum and subtractions between two monomials, and their final value is always a monomial. The numerical coefficients of monomials were sampled in the interval $[-99,99]$, and each monomial can contain up to four literal variables sampled from the set $\{a,b,x,y\}$. All monomials appearing in a given formula have the same literal variables. As in the case of Arithmetic, all intermediate numerical values were taken modulo $100$ when computing the final value of the formula.
@@ 5.2.1 Neural Rewriting System
We describe here how we built the training and validation sets for the Selector and Solver modules. We provide the details of the training procedure and hyperparameter selection for the models in the Technical Appendix.
Selector We have included in the training set for the Selector both top-level formulas and formulas that appear as intermediate solution steps in their simplification process. We did so in order to show the Selector during training the complete solution process of formulas up to a some level of difficulty. Indeed, we aim to learn a selection mechanism that can be applied also to more complex formulas.
In Arithmetic and Algebra problems we included in the training set formulas with nesting levels 1, 2 and 3, together with formulas that represent intermediate steps in the resolution process of these formulas. In the ListOps problem, in which we could specify the number of arguments that appear in the formulas, we included in the training set formulas with nesting levels 1 and 2 that have 2 or 3 arguments for each operator, together with formulas appearing in their resolution process. In all three problems, the training set also contains atomic elements which correspond to the final value of the initial formula. During training, we generated batches containing, with equal probability, samples with different structural complexity.
We constructed a separate in-distribution validation set with samples having the same structural characteristics as the those in training set. Since the Selector should possess out-of-distribution (OOD) generalization capabilities to identify the leaf formulas even in longer inputs than those seen during training, we also built a separate out-of-distribution validation set containing formulas with greater structural complexity. As done in the training set, we also included formulas representing examples of intermediate resolution steps. For Arithmetic and Algebra, we included samples with nesting levels 4, 5 and 6, together with formulas that appear in their resolution process. For ListOps, we included in the OOD validation set formulas with nesting levels 3 and 4 that have 2, 3 or 4 arguments for each operator, together with formulas that appear in their resolution process. In order to control the structural complexity of the formulas, the OOD validation sets are balanced across nesting level of the leaf expressions [^3] In the case of ListOps we also balance across the number of arguments in the leaf expression..
Solver The training and validation set of the Solver are composed by two kinds of samples: atomic elements and leaf formulas. During training, we generated batches including both kinds of samples with equal probability.
@@ 5.2.2 Neural Data Router
The Neural Data Router (NDR) is a variant of a transformer encoder designed to learn to solve algorithmic problems with strong out-of-distribution generalization capacity. This model has been previously evaluated on relatively simple algorithmic benchmarks that are nevertheless significantly similar to the formula simplification problems we consider, including the solution of formulas in the original ListOps dataset and simple arithmetic formulas on single-digit integers. The main differences with the problems we consider here is that the formulas that appear in our problems can have in general many leaf expressions and therefore can be simplified in more than one point, and the operands in our Arithmetic and Algebra problems have a greater complexity. Moreover, in the ListOps problems we aim to generalize on the number of operands as well as on the nesting depth of the formulas, whereas in the original work the number of operands in each expression was fixed. Furthermore, we use significantly less samples during training, and of lower complexity, as the NDR was originally trained on ListOps formulas with up to 5 nested operations and a variable number of operands between 2 and 5.
The problems considered in the original work always had a single-digit integer as a result, which the model was trained to output as the first token in the sequence given as output by the encoder. Since this no longer apply in our problems, except in the case of ListOps, in general we read the final answer from the first $k$ positions of the sequence produced by the encoder, where $k$ is the maximum length of a problem’s targets.
We build all development sets for the NDR using the same top-level formulas we included in the analogous sets for the Selector. Since the NDR was designed to compute intermediate results sub-symbolically, we use as training target the final simplified form of the input formula and we do not include intermediate formulas as we do with the Neural Rewriting System. Following the original experimental protocol, we made the training set balanced across nesting level and number of operands. As done when training the Selector module, we create in- and out-of-distribution validation sets to optimize hyperparameters, using the Weights and Biases Bayesian search on the same hyperparameters and intervals described in the original work. We report the final hyperparameter values for each task in the Technical Appendix.
@@ 5.2.3 GPT-4
Research on prompt engineering techniques has shown that appropriate prompting method can boost the reasoning capabilities of LLMs. We choose to prompt GPT-4 using self-consistency prompting Wang et al. (2023) combined with zero-shot CoT Kojima et al. (2022).
Zero-shot CoT has been introduced as a simpler alternative to Chain-of-Thought (CoT) prompting, which allows to achieve similar performance on reasoning benchmarks without the need to engineer exemplars for few-shot reasoning. This is achieved by simply making the answer of the model start with the sentence: “Let’s think step-by-step”. Self-consistency prompting is based on the idea that reasoning problems can often be solved following different reasoning paths that can lead to the same conclusion. Therefore, letting the model generate several outputs and then choosing the one which was generated more consistently can increase the confidence in generated output and improve accuracy. To this aim, we generate 10 outputs for each input, and indeed observe a significant improvement in the model’s performance. We report the performance of GPT-4 with zero-shot CoT
@@ 5.3 Results
We report the performance of all models considered in Figure 3. All models are tested on 100 samples for each level of complexity. We measure the sequence accuracy of the models, that is the exact match between the model’s output and the target sequence. When we evaluate GPT-4 on the Algebra task, we also take into account equivalent forms of the symbolic outputs using the SymPy library Meurer et al. (2017).
@@
We first observe that all the models have a better performance on simple data splits, which tends to degrade on more complex splits. This indicates that increasing the number of arguments and the nesting of formulas makes the problems more difficult for all models, as expected.
Analysing the performance of the Neural Data Router, we find that the models with the best hyperparameters configurations have a lower accuracy than other models on the simple splits and are not able to generalize the learned solution process to complex problem instances.
@@
We observe that GPT4 can achieve perfect or almost perfect results on the simpler problem instances, surpassing the performance of both the Neural Data Router and our modular architecture. However, the performance of GPT-4 is significantly lower on more complex instances of the problems. Indeed, while self-consistency prompting improves the model’s performance compared to vanilla zero-shot CoT prompting [^4] We report the performance of GPT-4 with vanilla zero-shot CoT prompting in the Technical Appendix for comparison., it has a limited impact on the capacity of the LLM to solve complex formulas. Comparing the performance achieved on ListOps to that on the other two tasks, we can conjecture that solving deeply nested formulas is harder for GPT-4 than solving formulas with many operands. Furthermore, it seems that accurately computing many arithmetic operations between double-digit integers, including multiplication modulo 100, is the hardest task for the LLM.
Figure 4: Errors committed by the Neural Rewriting System classified by cause of error. Error bars corresponding to a complexity split refer to models where the Selector generates 10, 100 or 1000 outputs per input, and employs the dynamic windowing mechanism. Complexity splits are defined by their nesting level N and number of arguments A.
@@
Figure 5: Input length against the average confidence score of 1000 outputs of the Selector. The vertical lines represent the maximum input length in the training set.
As for the Neural Rewriting System, generating 1000 outputs per input in the Selector allowed to achieve a higher degree of out-of-distribution generalization on more complex samples than those seen by the Selector during training. The performance of the model is consistently equal or superior to that of the Neural Data Router on both simple and complex data splits. Compared to GPT-4, the model achieves a similar performance on the simplest data splits, but its accuracy of is higher than that of GPT-4 on data splits of medium difficulty, such as arithmetic expressions of nesting level 4 and 5 and algebraic expression of nesting level 3 and 4. However, also the accuracy of the Neural Rewriting System drops dramatically on the most difficult arithmetic and algebraic formulas.
We analysed the root cause of the errors committed by the NRS model for each difficulty split in Figure 4. We also assessed the impact of the multi-output generation strategy on the model’s performance by comparing the errors committed by the model when 10, 100 and 1000 outputs per input were generated by the Selector. We classify the cause of the errors into three categories: errors caused by the Selector when it identifies a leaf formula which is not present in the input formula; errors caused by the Selector when it identifies a leaf formula that is present in the input formula but does not correspond to a well-formed expression and whose substitution thus causes the corruption of the input formula; and errors caused by the Solver when it outputs a wrong solution for the leaf formula identified by the Selector. We call these selector/missing, selector/corrupted and solver errors, respectively.
Looking at the frequency of each of these classes of errors, we clearly see that the main cause of error made on the hardest split is the Selector, either due to non-existing leaf formulas or ill-formed ones. However, we can also recognize that the multi-output generation strategy is effective for simpler data splits: generating many Selector outputs effectively reduces the amount of errors it causes, and especially the ones due to the identification of a non-existing leaf formula.
@@
Since the confidence score plays an important role in the selection of the best Selector output in the multi-output generation strategy, we analysed the variation of the score corresponding to variations of length of the input sequences in the test set. Figure 5 represents input lengths versus the average confidence score of 1000 Selector outputs. The vertical lines in the plot represent the maximum input length seen during training. While the average confidence score remains close to the training range for outputs that are moderately longer than the longest training samples, it drops significantly for sequences much longer than that. We also notice that confidence scores are lower for very short arithmetic and algebraic formulas. We suppose this is due to the fact that the number of formulas of a given nesting level grows exponentially with the value of nesting level. Therefore, the Selector observed fewer examples of formulas with nesting level 1 during training. However, this does not hinder performance when the Selector generates many outputs per input.
The dynamic windowing mechanism effectively addresses the drop of confidence on longer sequences. We report in Figure 3 the accuracy of a Neural Rewriting System in which the Selector generates 1000 outputs per input and applies the dynamic windowing mechanism on the longest sequences. We selected the length threshold $T$ that is used to specify which sequences will be processed using dynamic windowing based on the analysis of the confidence scores of the Selector discussed above. Specifically, we set the thresholds to 60, 150 and 200 for the ListOps, Arithmetic, and Algebra problems, respectively. The analysis of errors in Figure 4 shows that this mechanism improves the model’s performance by further reducing the number of errors caused by the Selector on the most complex data splits in all problems.

### v2

@@ 4.2 Solver Module
Unlike classical rewriting systems that use symbolic rule dictionaries, we learn rewriting rules directly from data by implementing the function $sol:F^{L}\rightarrow E$ with a trainable neural network. Given a leaf formula $f^{L}$ by the Selector, the Solver is trained to produce the equivalent reduction $e$ according to a valid rewriting rule. When training the Solver, we assume that the Selector always generates perfect outputs (i.e., we use teacher forcing). The Solver also learns to recognize the computation’s termination state, signaling when such a state is reached. It maps atomic elements, representing a formula’s final value, to the special symbol $\omega$, indicating the end of computation. Similar to the Selector, we frame the task as a sequence-to-sequence problem. We use a vanilla transformer encoder-decoder for the Solver since it learns simple input-output mappings.
@@ 4.3 Combiner Module
The last module in the architecture is the Combiner, a neural implementation of the function $com:F\times F^{L}\times E\rightarrow F$. Its purpose is thus to produce a simplified version of the original formula, given the formula itself $f$, the leaf formula $f^{L}$ identified by the Selector, and its reduction $e$ computed by the Solver.
@@
More precisely, we represent both the input sequence $f$ and the leaf formula $f^{L}$ as sequences of 1-hot vectors over the same vocabulary. Since the leaf formulas found for different sequences in a batch can have different lengths, we pad each one with zeros to prevent the padding to match in the input. Then, we set the filter of the 2D CNN to the 1-hot representation of $f^{L}$. We refer to the CNN parameterized in this way as $\mathrm{CNN}_{f^{L}}$. Doing so allows us to obtain from the output of the convolution both information on the location of the best match of $f^{L}$ in $f$ and on the number of tokens in $f^{L}$ that match $f$ exactly in some point. Indeed, we can compute the location of the best match as $\mathrm{pos}(f^{L},f)=\mathrm{argmax}~{}\mathrm{CNN}_{f^{L}}(f)$. Furthermore, we can compute the agreement score $a(f^{L},f)=\frac{\mathrm{max}~{}\mathrm{CNN}_{f^{L}}(f)}{|f^{L}|}$, where $|f^{L}|$ is the number of tokens in $f^{L}$. Dividing by $|f^{L}|$ makes the score normalized, which allows us to compare the agreement scores of leaf formulas with different lengths. Indeed, as described in Section 4, the Selector uses this score for multi-output generation to discard the outputs that do not have an exact match in the input formula. Notice that in this case the CNN is parameterized using candidate leaf formulas $\hat{f}^{L}_{j}$ whose accuracy scores with $f$ are compared. If there is no Selector output such that $a(\hat{f}^{L}_{j},f)=1$, the computation on the input sequence $f$ is stopped, and this is considered a failure of the model.
@@ 5 Experiments and results
Figure 3: Performance of the Neural Data Router, GPT-4 and Neural Rewriting System on ListOps, arithmetic and algebraic formulas.
We evaluated the capacity of our model to solve formula simplification problems on ListOps, arithmetic, and algebraic formulas. We compare the performance of the Neural Rewriting System to the Neural Data Router and GPT-4, one of the most advanced LLMs currently available. Code and data used in our experiments [25] and a Technical Appendix containing methodological details and additional results [27] are available.
@@ 5.1 Datasets
For all three tasks, the automatic generation of formulas is parameterized by specifying the nesting level of a formula. Any formula is nested at each level in two points: exactly two arguments in the formulas on that level will be other formulas. Formulas are tokenized at the character level for all tasks and trained models.
ListOps The ListOps dataset [22] was introduced to evaluate the capacity of neural networks to build parse trees of nested formulas. The original dataset consists of formulas composed of operations on lists of integers, including minimum, maximum, median and sum modulo $10$ of a list of integers. We modified the ListOps dataset so that formulas had exactly 2 nesting points at each nesting level, and we made it possible to specify the number of arguments that appear in formulas at any level. Since we are interested in the system’s capacity to generalize on highly nested formulas rather than learning specific operations, we reduced the set of operations to minimum, maximum and sum modulo $10$.
@@
Algebra We considered a subset of algebraic formulas that can always be reduced up to a minimal form in a deterministic way. Such formulas involve sum and subtractions between two monomials and their final value is always a monomial. The numerical coefficients of monomials were sampled in the interval $[-99,99]$, and each monomial can contain up to four literal variables sampled from the set $\{a,b,x,y\}$. All monomials appearing in a given formula have the same literal variables. As in the case of Arithmetic, all intermediate numerical values were taken modulo $100$ when computing the final value of the formula.
@@ 5.2.1 Neural Rewriting System
We describe here how we built the training and validation sets for the Selector and Solver modules. Further methodological details can be found in the Technical Appendix.
Selector In Arithmetic and Algebra problems we included in the training set formulas with nesting levels 1, 2 and 3. In the ListOps problem, in which we could specify the number of arguments that appear in the formulas, we included in the training set formulas with nesting levels 1 and 2 that have 2 or 3 arguments for each operator. In all three problems, the training set also contains atomic elements corresponding to the initial formula’s final value. Solving formula simplification problems by iteratively simplifying leaf formulas produces several simplifications of the initial formula. To show to the Selector the complete solution process of formulas in the training set, we have also included in it formulas that appear as intermediate solution steps of the formulas described above.
We constructed a separate in-distribution validation set with samples with the same structural characteristics as those in the training set. Unlike common machine learning tasks where models are tested on the same data distribution they are trained on, we want the Selector to possess out-of-distribution (OOD) generalization capabilities to identify the leaf formulas even in longer inputs than those seen during training. For this reason, we also built a separate out-of-distribution validation set containing formulas with greater structural complexity, and use this set for model selection. For Arithmetic and Algebra, we included samples with nesting levels 4, 5 and 6. For ListOps, we included formulas with nesting levels 3 and 4 with 2, 3 or 4 arguments for each operator. To select the most capable model across the iterative resolution process, we also included formulas representing examples of intermediate resolution steps. To control the structural complexity of the formulas, the OOD validation sets are balanced across the nesting level of the leaf expressions. [^3] In the case of ListOps, we also balance across the number of arguments in the leaf expression.
Solver The training and validation set of the Solver included two kinds of samples: leaf formulas, mapped to equivalent atomic elements, and atomic elements themselves, mapped to the end-of-computation special symbol $\omega$. During training, we generated batches that included both kinds of samples with equal probability to avoid biasing the model toward solving leaf formulas.
@@ 5.2.2 Neural Data Router
The Neural Data Router (NDR) is a variant of a transformer encoder designed to learn to solve algorithmic problems with strong out-of-distribution generalization capacity. In terms of the formalization of formula simplification problems we have defined in Section 3, the model takes as input a formula $f$ and gives as output the equivalent atomic element $e$. The NDR thus implements the composition of the $sel$, $sol$, and $com$ functions in one or more transformer layers, assuming that the effective integration between the three happens sub-symbolically in the embedding space during training.
This model has been previously evaluated on relatively simple algorithmic benchmarks that are nevertheless significantly similar to the problems we consider, including the solution of formulas in the original ListOps dataset and simple arithmetic formulas on single-digit integers. The main difference with the problems we consider here is that the operands in our Arithmetic and Algebra problems are of greater complexity. Moreover, in the ListOps problems, we aim to generalize both on the number of operands and on the nesting depth of the formulas, whereas, in the original work, the number of operands observed at training time was the same as that observed at test time. Furthermore, we use significantly fewer samples during training and of lower complexity, as the NDR was originally trained on arithmetic and ListOps formulas with up to 5 nested operations.
We applied a minor modification to the architecture to adapt the model to the problems we have considered. Indeed, the problems considered in the original work always had a single-digit integer as a result, which the model was trained to output as the first token in the sequence given as output by the encoder. Since, in general, this no longer applies to our problems, we read the final answer from the first $k$ positions of the sequence produced by the encoder, where $k$ is the maximum length of a problem’s targets.
We build all development sets for the NDR using the same top-level formulas we included in the analogous sets for the Selector. Since the NDR was designed to compute intermediate results sub-symbolically, we use as training target the final simplified form of the input formula and we do not include intermediate formulas as we do with the Neural Rewriting System.  [^4] For comparison, we include in the Technical Appendix results for models trained on larger datasets that also include intermediate formulas. Following the original experimental protocol, we made the training set balanced across nesting levels and number of operands. As done when training the Selector module, we create in- and out-of-distribution validation sets and use the latter to optimize hyperparameters, using the Weights and Biases Bayesian search on the same hyperparameters and intervals described in the original work. We report the final hyperparameter values for each task in the Technical Appendix.
@@ 5.2.3 GPT-4
Chain-of-Though (CoT) prompting improves the performance of LLMs on reasoning tasks by enabling step-by-step solution procedures. In terms of the formalization outlined in Section 3, we may observe how such reasoning chains resemble the iterative solution method of nested formulas based on rewriting. Therefore, we can say that LLMs prompted using CoT-like methods loosely implement the composition of $sel$, $sol$, and $com$ functions in the whole model, where the iterative application of the function is implemented via the auto-regressive generation of outputs. [^5] Notice that in general, at any given point of the solution process, the context given as input of the model will contain several different simplifications of the input function $f$ rather than just one as it happens with the other models we consider.
We choose to prompt GPT-4 using self-consistency prompting [32] combined with zero-shot Chain-of-Thought (CoT) [12]. Zero-shot CoT has been introduced as a simpler alternative to CoT prompting, which allows users to achieve similar performance on reasoning benchmarks without the need to engineer exemplars for few-shot reasoning. This is achieved by simply making the model’s answer start with the sentence: “Let’s think step-by-step”. After the model has generated a response, it is prompted again to retrieve a well-formatted output. Self-consistency prompting operates on the concept that reasoning problems can have multiple valid paths leading to the same conclusion. To harness this, we generate 10 outputs for each input, allowing us to select the most consistently produced one. This method boosts confidence in the model’s output and significantly enhances accuracy, resulting in a notable improvement in performance.
@@ 5.3 Results
We report in Figure 3 the performances of all considered models. All models are tested on 100 samples for each level of complexity. We measure the sequence accuracy of the models, that is the exact match between the model’s output and the target sequence. When we evaluate GPT-4 on the Algebra task, we take into account equivalent forms of the symbolic outputs using the SymPy library [20].
@@
We first observe that all the models perform better on simple data splits and tend to worsen on more complex ones. This indicates that increasing the number of arguments and the nesting of formulas makes the problems more difficult for all models, as expected.
Analyzing the performance of the Neural Data Router, we find that the models with the best hyperparameters configurations have lower accuracy than other models on the simple splits and are not able to generalize the learned solution process to complex problem instances. Learning to solve simple formulas with a single nesting level – i.e., learning the $sol$ function – appears to be the easiest task for the model, although it is learned with almost perfect accuracy only on arithmetic formulas. The lack of generalization capacity on complex formulas may, therefore, be determined by two different situations: making one or more errors while solving many leaf formulas, or learning a composition of the $sel$, $sol$, and $com$ functions that cannot generalize effectively on nested formulas. Since the model is unable to generalize both on arithmetic and algebraic formulas, whose leaf formulas can be solved with substantially different levels of accuracy, we argue that both factors could play a role in causing the limited generalization capacity of the model.
@@
We observe that GPT-4 can achieve perfect or almost perfect results on leaf formulas and simpler problem instances, surpassing the performance of the Neural Data Router and, in some cases, our architecture. However, the performance of GPT-4 is significantly lower on deeply nested formulas. Indeed, while self-consistency prompting improved the model’s performance compared to vanilla zero-shot CoT prompting [^6] We report the performance of GPT-4 with vanilla zero-shot CoT prompting in the Technical Appendix for comparison., it has a limited impact on the capacity of the LLM to solve complex formulas, consistently with previous research [26].
Considering the model as an approximation of the composition of the $sel$, $sol$, and $com$ functions, its limited generalization capability seems to be determined by the poor quality of this approximation obtained with the prompting method we used. However, by comparing the performance of the model on ListOps formulas with the other two tasks, we observe that the complexity of the individual leaf formulas (much simpler in the case of ListOps) seems to influence the capability of the model to apply a solution method that can generalize systematically on complex formulas.
@@
Figure 4: Input length against the average confidence score of 1,000 outputs of the Selector. The vertical lines represent the maximum input length in the training set.
Coherently with the model selection process, in Figure 3, we report the performance of our Neural Rewriting System with 1,000-way multi-output generation, and value of the Dynamic Windowing threshold $T$ of 60, 150 and 200 for the ListOps, Arithmetic and Algebra problems, respectively. The Neural Rewriting System consistently outperforms the baselines on the three formula simplification tasks we consider. Importantly, the system shows a significantly high degree of generalization on formulas that are much more complex than those observed during training. Given the modular nature of our architecture, in the following paragraphs, we will analyze the role of each architecture component in achieving this result.
Figure 5: Errors committed by the Neural Rewriting System classified by cause of error. Error bars corresponding to a complexity split refer to models where the Selector generates 10, 100 or 1,000 outputs per input and (optionally) employs the Dynamic Windowing mechanism. Data splits are defined by their nesting level N and number of arguments A.
@@
We first analyze the variation of the confidence score corresponding to variations in the length of the input sequences in the test set. We do so because the confidence score plays an important role in selecting the best Selector output in multi-output generation. Figure 4 represents input lengths versus the average confidence score of 1,000 Selector outputs. The vertical lines in the plot represent the maximum input length seen during training. While the average confidence score remains close to the training range for outputs that are moderately longer than the longest training samples, it drops significantly for sequences much longer than that. This can directly determine a decrease in model performance on very long formulas since Selector outputs with a low confidence score are more likely to be incorrect. Indeed, we set the value of the Dynamic Windowing threshold $T$, which determines the sequences on which the mechanism will be applied, by looking at the points where Selector confidence drops in each task. We also notice that confidence scores are lower for very short arithmetic and algebraic formulas. We suppose this is due to the fact that the number of formulas of a given nesting level grows exponentially with the value of the nesting level. Therefore, the Selector observed fewer examples of formulas with nesting level 1 during training. However, this does not hinder performance when the Selector generates many outputs per input.
We finally assessed the impact of the multi-output generation strategy on the model’s performance by comparing the errors committed by the model when the Selector generated 10, 100 and 1,000 outputs per input. We measure the errors separately on subsets of formulas with different nesting levels and number of operands, to understand the impact of increasing the number of outputs generated by the model on the solution of both simple and complex formulas. Furthermore, for each model configuration and each subset of samples, we analyzed the root cause of the errors committed by the NRS. We classify the cause of the errors into three categories: errors caused by the Selector when it outputs a leaf formula that is not present in the input formula (missing-leaf); errors caused by the Selector when it outputs a leaf formula that is present in the input formula but does not correspond to a well-formed expression and whose substitution thus causes the corruption of the input formula (corrupted-leaf); and errors caused by the Solver when it outputs a wrong solution for the leaf formula identified by the Selector (wrong-solution). Figure 5 represents this multi-faceted measurement of the errors committed by the models.
Looking at the frequency of each of these classes of errors, we clearly see that generating only 10 outputs per input with the Selector is particularly ineffective on algebraic formulas, and on deeply nested arithmetic and ListOps formulas. Increasing the number of generated outputs to 100 or 1,000 considerably improves the System’s performance on most formulas of all three tasks, apart from the most deeply nested arithmetic and algebraic formulas. On such samples, applying Dynamic Windowing drastically reduces the number of errors the model commits while also bringing further minor improvements on simpler formulas.
As we described in Section 4, the modules in the Neural Rewriting System have been designed as neural implementations of the $sel$, $sol$ and $com$ functions that appear in the simple solution algorithm for formula simplification problems described in Section 3. For this reason, we can analyze in a greater level of detail to which modules are most error due, and thus which of the functions listed above have been more difficult to reproduce in a neural network.
Given our design choice to implement the $com$ function using a CNN dynamically parameterized by the Selector, the errors due to this module can be directly traced back to errors made by the Selector. Looking at the classification of errors represented in Figure 5, we can see how the errors due to a wrong-solution of a leaf formula by the Solver are the minority. Therefore, the $sol$ function appears to be relatively easy to implement in a transformer, consistent with our prior observations on NDR and GPT-4 models. Errors due to the Selector are instead the vast majority and are only partially mitigated by the generation of multiple outputs and the Dynamic Windowing, especially on algebraic formulas. The $sel$ function, which formalizes the ability to identify leaf expressions, appears to be the hardest one to model with a transformer-based architecture, particularly due to the necessity to generalize on very long formulas.